OpenStudy (anonymous):

The following system is solved by substitution. Which expression is substituted for z in the second and third equations? x + 5y – z = 6 3x – y – 2z = 4 3x – 2y + 3z = 13

OpenStudy (chaise):

x+5y-z=6 If we rearrange this to make z the subject we can substitute this into the last two equations. x+5y-6=z z=x+5y-6

OpenStudy (anonymous):

Why thank you again :) lol